### v3

# Reconfigurable Intelligent Surface-Assisted Localization in OFDM Systems with Carrier Frequency Offset and Phase Noise

Abstract
Reconfigurable intelligent surface (RIS)-assisted communication systems have been extensively studied for providing high-precision location services. However, most studies have overlooked the impact of carrier frequency offset (CFO) and phase noise (PN) resulting from hardware impairments on localization. This paper presents a novel, alternating optimization (AO)-based algorithm to jointly estimate the CFO, PN, and user position, where, provided the user position, closed-form expressions for the CFO and PN are derived per iteration, significantly reducing the complexity and enhancing the stability of the algorithm. Another important aspect is a new RIS phase shift optimization algorithm developed to minimize the analytical lower bound of localization accuracy, hence benefiting localization. The semidefinite relaxation method and Schur complement are utilized to convexify this challenging non-convex optimization problem to a semidefinite program. Simulations demonstrate the effectiveness of the proposed algorithms, with the localization accuracy enhanced by two orders of magnitude. The localization accuracy of the proposed algorithm is close to the analytical lower bound, with a root mean square error of lower than $\rm 10^{-2}\>m$.
Index Terms:
Localization, reconfigurable intelligent surface, phase shift configuration, carrier frequency offset, phase noise.
I Introduction
I-A Motivation and Challenges
With the development of industrial Internet of Things, there is an increasing demand for centimeter-level localization accuracy in the sixth-generation (6G) mobile networks [1, 2, 3]. While global navigation satellite systems (GNSSs) have been widely applied to outdoor scenarios [4], they undergo insufficient localization accuracy and signal instability arising from obstacles [5, 6]. In recent years, communication systems have been considered to assist in localization, e.g., integrated sensing and communication (ISAC), aiming to achieve more accurate and continuous localization alongside GNSS signals.
Although communication-assisted localization systems have been extensively studied, few studies have considered the adverse effects of hardware impairments, such as carrier frequency offset (CFO) and phase noise (PN), on localization accuracy. CFO is a constant parameter due to the frequency offset between the transmitter and the receiver. Doppler shift can also lead to CFO. PN is a random process resulting from the fluctuation of the oscillators at the transmitters and receivers [7]. CFO and PN can result in signal constellation rotations in single carrier communication systems, or inter-carrier interference (ICI) in broadband communication systems, for example, orthogonal frequency division multiplexing (OFDM) systems. The overall communication performance and the localization accuracy can degrade significantly, due to the CFO and PN.
Another challenge arises from the increasingly likely blockages of signals, e.g., millimeter wave (mmWave) signals with the quasi-optical property, in many dense urban areas [8]. Reconfigurable intelligent surface (RIS) [9, 10] has been considered in communication systems to deal with the problem of signal blockage. Localization can be implemented with the help of virtual line-of-sight (LoS) paths provided by RISs even when the LoS paths are blocked, and can be improved via passive beamforming. In this sense, RIS-assisted user localization is of practical interest, and could be particularly challenging with the further consideration of the CFO and PN. However, this has not been studied in the existing literature.
I-B Related works
Communication systems have been widely employed to assist in high-precision user localization, typically under a far-field setting; i.e., the user and the anchors are in the far field of each other. For example, Shahmansoori et al. [11] proposed a three-stage localization algorithm that improves the simultaneous orthogonal matching pursuit algorithm and the space alternating generalized expectation-maximization algorithm, and exploited the sparsity of millimeter-wave multiple-input multiple-output (MIMO) channels to simplify the channel model. Hu et al. [12] presented a two-dimensional (2D) localization algorithm based on the celebrated estimating signal parameters via rotational invariance technique (ESPRIT), which estimates user positions by constructing signal subspaces.
Some studies [13, 14] have explored the impact of realistic hardware impairments, including the CFO and PN, with no RIS considered. For instance, Meles et al. [13] simulated the effects of CFO and PN on angle-of-arrival (AOA) estimation. Chen et al. [14] considered the impacts of CFO, PN, mutual coupling, and power amplifier nonlinearity on 6G localization, and performed analysis using the derived misspecified Cramér-Rao lower bound (CRLB). However, neither of them provided an estimation algorithm to handle the hardware impairments.
Some other works [15, 16, 17, 18] have delved into localization algorithm design for scenarios with either CFO or PN. On the one hand, the CFO was studied in [15, 16]. Qiu et al. [15] designed an maximum likelihood estimation (MLE)-based weighted iterative algorithm for the joint estimation of CFO and AOA, yet the coupling of CFO and position parameters in each iteration incurs high complexity. Zakariyya et al. [16] proposed a high-resolution multiple signal classification (MUSIC) algorithm for localization, but did not utilize the CFO estimates in localization, leading to a loss in localization accuracy. On the other hand, the PN was considered in [17, 18]. Xiao et al. [17] combined the expectation maximization (EM) and MUSIC algorithms for joint estimation of AOA and PN, but only considered a single-carrier system under a 2D setting. Zheng et al. [18] formulated a maximum a-posteriori (MAP) problem iteratively solved using the block majorization minimization (MM) algorithm after approximation, where the PN was simplified as a Gaussian variable. However, the above methods have not considered the use of RISs nor addressed the signal blockage by obstacles.
RIS has been integrated into localization systems to address the problem of signal blockage, but little consideration has been given to the above-mentioned hardware impairments, i.e., CFO and PN. From the perspective of user mobility, existing studies can be divided between static localization [19, 20, 21] and dynamic localization [22, 23]. The former encompasses scenarios where the users undergo minimal movement within a coherent time. The latter involves scenarios when the users move at a relatively high speed, and the Doppler shift cannot be overlooked. From the perspective of localization distance, existing research can be divided between far-field localization [21, 23, 24] and near-field localization [20, 21, 23, 25]. In the far-field case, the user-RIS distance exceeds Fraunhofer distance [26], and the signal can be approximated as a planar wave. Conversely, in the near-field case, the signal cannot be treated as parallel propagation and should be considered as a spherical wave.
For example, Luan et al. [27] focused on the near-field static 2D localization scenario, and proposed a robust optimization algorithm to configure the RIS phase shifts for improving localization accuracy. The transmit powers were also optimized in the considered localization system. Ozturk et al. [28] further introduced a more practical RIS model with phase-dependent amplitude variations in the near-field circumstance. The misspecified CRLB was derived and an approximated mismatched maximum likelihood-based algorithm was proposed for joint UE localization and RIS parameter estimation. The system model in this paper also belongs to the near-field static category. Although we adopt the classical ideal RIS phase shift model with unit amplitude for simplicity, we first take the CFO and PN into consideration. Elzanaty et al. [23] studied RIS-assisted localization for multi-antenna mobile users. They proposed a three-dimensional (3D) model that applies to both far-field and near-field scenarios, and analyzed the position and orientation error bounds. Lin et al. [29] introduced a novel twin-RIS configuration comprising two RISs with relative spatial rotations. To extract the parameters of the cascaded channel, the authors employed array signal processing and atomic norm denoising techniques. They established nonlinear equations exploiting the characteristics of the twin-RIS structure to recover the multipath parameters. User localization for static users in the far-field scenario was then performed based on the estimated channel parameters. Teng et al. [22] introduced a new message-passing algorithm called BULT to locate mobile users in the far-field scenario. The Bayesian CRLB was derived for BULT.
The existing RIS-assisted localization studies have assumed ideal channel conditions without considering realistic hardware impairments. Only several studies have considered modeling some specific impairments. For example, the studies in [23, 24, 30] took into account clock offset, and the work in [30] considered phase offsets. Zhang et al. [31] considered the CFO while overlooking the effects of PN. They proposed a MUSIC-based algorithm that directly estimates the position parameters together with the CFO. However, the algorithm is only suitable for far-field scenarios. To the best of our knowledge, no existing research has specifically analyzed the impact of both CFO and PN on RIS-assisted near-field localization systems.
I-C Contributions
In this paper, we investigate a RIS-assisted OFDM system, where practical CFO and PN convolve with the data symbols, causing signal constellation rotation and ICI and, consequently, hindering user localization in near-field environments. The key contributions of this paper are outlined as follows:
  • We investigate a problem of static near-field localization in a RIS-aided OFDM system, while taking into account the CFO and PN. We propose an alternating optimization (AO)-based iterative estimation approach, where the MAP criterion and gradient descent (GD) algorithm are employed with closed-form expressions derived for the CFO and PN estimation.
  • We derive the hybrid CRLB (HCRLB) for the considered joint estimation problem. The phase ambiguity in CFO and PN estimation is analyzed. To evaluate the accuracy of joint CFO and PN estimation, a joint mean square error (MSE) metric is introduced.
  • We further optimize the phase shifts of the RIS to improve the performance of localization. By applying the semi-definite relaxation (SDR) and the Schur complement, we convert the non-convex optimization problem into a convex semi-definite programming (SDP) problem solved efficiently.
  • We perform comprehensive numerical simulations to assess the performance of the proposed algorithms. It is shown that the localization accuracy is improved by two orders of magnitude compared to random RIS phase shift configuration. Compared to the state-of-the-art benchmarks, our algorithms can improve the localization accuracy by about one order of magnitude under relatively low transmit power and even by two orders of magnitude under high transmit power.
Figure 1: The considered RIS-aided 3D near-field localization system, including an AN, a UE, and an $N_{R}$-element RIS. The LoS link between the AN and UE is obstructed. The UE utilizes the signals transmitted by the AN and reflected by the RIS to localize its position.
The rest of this paper is organized as follows. Section II presents the system model. Section III presents the joint estimation algorithm. Section IV derives the HCRLB for the joint estimation algorithm. The proposed RIS phase shift optimization method is delineated in Section V. Section VI showcases comprehensive simulation results, and the paper is concluded in Section VII.
Notation: $\mathbb{E}(\cdot)$ and $p(\cdot)$ denote expectation and probability density function (PDF), respectively. ${\rm tr}(\cdot)$, $\det(\cdot)$, $(\cdot)^{\top}$, $(\cdot)^{*}$, $(\cdot)^{\rm H}$, and $(\cdot)^{\rm-1}$ denote trace, determinant, transpose, conjugate, hermitian, and inverse, respectively. ${\rm Diag}(\bm{v})$ indicates a diagonal matrix where the entries of vector $\bm{v}$ appear on its main diagonal. $\bm{v}[n]$ denotes the $n$-th entry of vector $\bm{v}$. ${\rm Blkdiag}(\mathbf{M}_{0},\mathbf{M}_{1},\cdots,\mathbf{M}_{N-1})$ denotes a block diagonal matrix with $\mathbf{M}_{0},\mathbf{M}_{1},\cdots,\mathbf{M}_{N-1}$ being the matrices along its main diagonal. $\left[\mathbf{M}\right]_{i:j,:}$ refers to a sub-matrix encompassing rows $i$ to $j$ of matrix $\mathbf{M}$, while $\left[\mathbf{M}\right]_{:,i:j}$ denotes a sub-matrix encompassing columns $i$ to $j$ of matrix $\mathbf{M}$. $\bm{0}_{M\times N}$ and $\mathbf{I}_{M}$ denote the $M\times N$ all-zero matrix and the $M\times M$ identity matrix, respectively. ${}\Re\left\{\cdot\right\}$ and ${}\Im\left\{\cdot\right\}$ represent the real and imaginary components of a complex value, respectively. $\mathbb{C}^{m\times n}$ and $\mathbb{R}^{m\times n}$ collect all $m\times n$ matrices with complex and real entries, respectively. $\mathcal{CN}(\bm{\mu},\mathbf{\Sigma})$ and $\mathcal{N}(\bm{\mu},\mathbf{\Sigma})$ represent complex and real Gaussian distributions, respectively, with a mean vector $\bm{\mu}$ and covariance matrix $\mathbf{\Sigma}$. Unless otherwise specified, all indices in the paper start from 0. Moreover, we summarize the notation used in this paper in Table I.
TABLE I: Notation and definition.
Notation Definition
$N_{R}$ Number of RIS elements
$N$ Number of subcarriers
$L$ Length of the approximate PN vector
$d$ Element spacing of the RIS
$D$ Aperture of the RIS
$f_{c}$ Carrier frequency of the system
$B$ Bandwidth of the system
$\lambda$ Wavelength of the system
$c$ Speed of light
$P$ Transmit power at the AN
$\Delta f$ CFO
$\phi$ Normalized CFO
$U$ Position sampling amount
$\sigma^{2}$ Variance of the noise at the UE
$\sigma_{\Delta}^{2}$ Innovation variance for PN
$\bm{p}_{a}$ Coordinate of the AN
$\bm{p}_{r}$ Coordinate of the reference point of the RIS
$\bm{p}_{u}$ Coordinate of the UE position
$\bm{\theta}$ PN vector
$\bm{\eta}$ Approximate PN vector
$\bm{y}$ Received signal at the UE
$\bm{\mu}$ Noise-free received signal at UE
$\bm{h}$ AN-RIS-UE frequency-domain channel
$\bm{w}$ RIS phase shift vector
$\mathbf{s}$ Transmitted symbol sequence at the AN
$\bm{v}$ Additive noise at UE
$\bm{\lambda}$ Parameter vector to be estimated
$\bm{\alpha}$ Transformed parameter vector
$\mathbf{B}$ Bayesian information matrix
$\bm{\Lambda}_{\phi}$ CFO matrix
$\bm{\Lambda}_{\bm{\theta}}$ PN matrix
$\mathbf{\Pi}$ Transition matrix for the PN vector
$\mathbf{\Xi}$ Transition matrix for the HCRLB
II System Model
The considered RIS-assisted OFDM downlink wireless localization system is composed of an anchor node (AN), a user equipment (UE), and an $N_{R}$-element RIS, as illustrated in Fig. 1. We make the assumption that the LoS path between the AN and UE is obstructed, and the channel state information (CSI) remains unchanged during localization, as in [21, 27, 30]. The positions of the AN and each RIS element, denoted as $\bm{p}_{a}$ and $\bm{p}_{r}=[x_{r},y_{r},z_{r}]^{\top}$, respectively, are known, where $r$ is the index of a RIS element. The position of UE, denoted as $\bm{p}_{u}=[x_{u},y_{u},z_{u}]^{\top}$, is to be estimated.
The RIS elements are organized in a uniform planar array (UPA). We take the bottom-left corner of the RIS as the reference point and place it at the coordinate origin. The RIS is placed on the $yz$-plane, and both edges on the coordinate axes are located in the positive half-axis. Hence, the coordinates of the $r$-th RIS element can be represented as $\bm{p}_{r}=[0,d\times((r-1)\mod N_{R}),d\times\lceil\frac{r}{\sqrt{N_{R}}}-1\rceil]^{\top}$, where $d$ stands for the element spacing of the RIS, and is typically half of a wavelength. To satisfy the conditions of the near-field setup, both the AN-RIS distance, as well as the RIS-UE distance, should lie within the Fresnel region [26], i.e., the interval $[0.62\sqrt{D^{3}/\lambda},2D^{2}/\lambda]$, where $D=\sqrt{2}d(\sqrt{N_{R}}-1)$ represents the aperture of the RIS and $\lambda$ denotes the wavelength.
Additionally, we adopt a polar coordinate system for estimating the user’s position, i.e., using $d_{Ru}$, $\phi_{az}$ and $\phi_{el}$, to express the user’s position, indicating the distance, azimuth angle, and elevation angle of UE relative to the coordinate origin, respectively. The relationships between these three parameters and $x_{u}$, $y_{u}$ and $z_{u}$ are given by
$d_{Ru}$ $=\sqrt{x_{u}^{2}+y_{u}^{2}+z_{u}^{2}};$ (1a)
$\phi_{az}$ $=\arctan\frac{y_{u}}{x_{u}};$ (1b)
$\phi_{el}$ $=\arccos\frac{z_{u}}{\sqrt{x_{u}^{2}+y_{u}^{2}+z_{u}^{2}}},$ (1c)
and, in turn, $x_{u}$, $y_{u}$ and $z_{u}$ can be written as
$x_{u}$ $=d_{Ru}\sin\phi_{el}\cos\phi_{az};$ (2a)
$y_{u}$ $=d_{Ru}\sin\phi_{el}\sin\phi_{az};$ (2b)
$z_{u}$ $=d_{Ru}\cos\phi_{el}.$ (2c)
An OFDM symbol with a center carrier frequency of $f_{c}$, a signal bandwidth of $B$ and $N$ subcarriers is considered in this paper. Upon sampling at a rate of $T_{s}=\frac{1}{NB}$ and removing the cyclic prefix (CP) at the UE, the received time-domain signal $\bm{y}$ can be written as
$\bm{y}=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\bm{h}_{LoS}+\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\bm{h}_{NLoS}+\tilde{\bm{v}},$ (3)
where
  • $P$ is the transmit power;
  • $\bm{\Lambda}_{\phi}$ is the CFO matrix resulting from a frequency mismatch between the transmitter and receiver. $\bm{\Lambda}_{\phi}$ is a diagonal matrix with $\left[\bm{\Lambda}_{\phi}\right]_{n,n}=e^{j2\pi\frac{n\Delta f}{NB}}=e^{j2\pi\frac{n}{N}\phi}$ being the $(n,n)$-th element along its main diagonal, where $\Delta f$ is the CFO, and $\phi=\frac{\Delta f}{B}$ is called the normalized CFO;
  • $\bm{\Lambda}_{\bm{\theta}}\triangleq{\rm Diag}\left[e^{j\bm{\theta}[0]},e^{j\bm{\theta}[1]},\cdots,e^{j\bm{\theta}[N-1]}\right]$ is the PN matrix. Drawing upon the attributes of the PN observed in practical oscillators, the PN vector $\bm{\theta}$ can be modeled as a Wiener process [32], i.e.,
    $\bm{\theta}[n]=\bm{\theta}[n-1]+\bm{\Delta}[n],$ (4)
    where $\bm{\Delta}[n]$ is a Gaussian variable, i.e., $\bm{\Delta}[n]\sim\mathcal{N}(0,\sigma_{\Delta}^{2})$. As in [33], we assume that $\bm{\theta}[-1]=0$. Then, the PN vector $\bm{\theta}\triangleq\left[\bm{\theta}[0],\bm{\theta}[1],\cdots,\bm{\theta}[N-1]\right]^{\top}$ yields $\bm{\theta}\sim\mathcal{N}(\bm{0}_{N\times 1},\bm{\Psi})$, where $\mathbf{\Psi}$ is given by
    $\mathbf{\Psi}=\sigma_{\Delta}^{2}\begin{bmatrix}1&1&1&\cdots&1&1\\ 1&2&2&\cdots&2&2\\ 1&2&3&3&\cdots&3\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 1&2&3&\cdots&N-1&N\end{bmatrix};$ (5)
  • $\mathbf{S}\triangleq\sqrt{N}\mathbf{F}^{\rm H}{\rm Diag}(\bm{s})$, where $\mathbf{F}^{\rm H}$ is the inverse fast Fourier transform (FFT) matrix and $\bm{s}$ is the frequency-domain transmitted symbol sequence at the AN;
  • $\bm{h}_{LoS}$ and $\bm{h}_{NLoS}$ represent the LoS and non-LoS parts of the frequency-domain channel, respectively;
  • $\tilde{\bm{v}}$ is the additive Gaussian noise at the UE, and follows the distribution $\tilde{\bm{v}}\sim\mathcal{CN}(\bm{0}_{N\times 1},\sigma_{AWGN}^{2}\mathbf{I}_{N})$.
Assume that $\bm{h}_{NLoS}$ follows a known Gaussian distribution $\bm{h}_{NLoS}\sim\mathcal{CN}(0,\sigma_{NLoS}^{2}\mathbf{I}_{N})$. Thus, $\bm{\mu}_{NLoS}\triangleq\mathbf{X}\bm{h}_{NLoS}$ also follows a known Gaussian distribution $\bm{\mu}_{NLoS}\sim\mathcal{CN}(0,\sigma_{NLoS}^{2}\mathbf{X}\mathbf{X}^{\rm H})$ with $\mathbf{X}\triangleq\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}$. Combine $\bm{\mu}_{NLoS}$ with the additive Gaussian noise $\tilde{\bm{v}}$. For conciseness, denote $\bm{h}_{LoS}$ with $\bm{h}$. The received signal is rewritten as
$\bm{y}=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\bm{h}+\bm{v},$ (6)
where $\bm{v}\triangleq\bm{\mu}_{NLoS}+\tilde{\bm{v}}$ is Gaussian distributed, i.e, $\bm{v}\sim\mathcal{CN}(0,\sigma^{2}\mathbf{I}_{N})$, with $\sigma^{2}=\sigma_{AWGN}^{2}+\sigma_{NLoS}^{2}$.
Each element of $\bm{h}$ can be written as $\bm{h}[n]=\sum_{r=1}^{N_{R}}\sqrt{\rho_{aR,n,r}}\sqrt{\rho_{Ru,n,r}}\bm{w}[r]e^{-j2\pi\frac{nB}{N}(\tau_{ar}+\tau_{ru})}$, where $\sqrt{\rho_{aR,n,r}}=\frac{\sqrt{G_{a}}\lambda_{n}}{4\pi d_{ar}}$ and $\sqrt{\rho_{Ru,n,r}}=\frac{\sqrt{G_{t}}\lambda_{n}}{4\pi d_{ru}}$ are the free-space path losses from the AN to the $r$-th RIS element, and from the $r$-th RIS element to the UE, respectively. $\bm{w}[r]$ is the phase shift of the $r$-th RIS element, $\tau_{ar}=\frac{d_{ar}}{c}=\frac{\|\bm{p}_{r}-\bm{p}_{a}\|}{c}$ is the delay from the AN to the $r$-th RIS element, while $\tau_{ru}=\frac{d_{ru}}{c}=\frac{\|\bm{p}_{u}-\bm{p}_{r}\|}{c}$ denotes the delay from the $r$-th RIS element to the UE, with the light speed $c$.
III Proposed Joint CFO, PN and UE Position Estimation Algorithm
In previous works [27, 30], localization was based on frequency-domain signals, and different subcarriers’ signals were often assumed to be independent. However, the CFO and PN would result in ICI, which increases the difficulty in the estimation of the CFO, PN and the position of the UE. The estimation can only be conducted in the time domain. In this section, we propose an AO-based iterative algorithm, which employs the MAP criterion and GD algorithm to jointly estimate the normalized CFO, PN, and the UE position.
Define the parameter vector $\bm{\lambda}\triangleq\left[\phi,\bm{\theta}^{\top},d_{Ru},\phi_{az},\phi_{el}\right]^{\top}$, wherein $\bm{\theta}$ is a random variable and the others are deterministic. The posterior distribution of $\bm{\lambda}$, given the received signal $\bm{y}$, can be expressed as
$p(\bm{\lambda}\mid\bm{y})=\frac{p(\bm{y}\mid\bm{\lambda})\times p(\bm{\theta})}{p(\bm{y})}.$ (7)
According to the MAP criterion, $\bm{\lambda}$ can be estimated by optimizing the following unconstrained function:
$\hat{\bm{\lambda}}\propto\arg\min_{\bm{\lambda}}\mathcal{L}(\bm{\lambda}),$ (8)
where $\mathcal{L}(\bm{\lambda})=-\log p(\bm{y}\mid\bm{\lambda})-\log p(\bm{\theta})$, which is obtained by calculating the negative logarithmic function of (7) and then suppressing constants.
III-A Phase Noise Estimation
III-A1 Approximation of PN vector
Figure 2: Eigenvalues of $\bm{\Psi}$ at the number of subcarriers $N=32$, and the innovation variance for PN $\sigma_{\Delta}^{2}=10^{-3}$.
From the definition of $\bm{\lambda}$, many variables need to be jointly estimated at the UE. To lower the computational complexity of the estimation, we exploit the redundancy of the covariance matrix of the PN vector to approximate it and reduce variables requiring estimation.
As illustrated in Fig. 2, the eigenvalues of $\bm{\Psi}$ are mostly close to zero, indicating significant redundancy in $\bm{\Psi}$. With this property, we can approximate the PN vector as
$\bm{\theta}\approx\bm{\Pi}\bm{\eta},$ (9)
where $\bm{\eta}\sim\mathcal{N}(\bm{0}_{L\times 1},\mathbf{I}_{L})$ contains $L\leq N$ PN variables, and $\bm{\Pi}\in\mathbb{R}^{N\times L}$ is the transition matrix. We set $L=\frac{N}{2}$ in this paper. To derive $\bm{\Pi}$, we first perform eigenvalue decomposition (EVD) on $\bm{\Psi}$, i.e., $\bm{\Psi}=\mathbf{U}\mathbf{D}\mathbf{U}^{\top}$, where $\mathbf{D}={\rm Diag}(\bm{e})$ and $\bm{e}$ collects the eigenvalues of $\bm{\Psi}$ arranged in descending order. Subsequently, $\bm{\Pi}$ can be calculated by $\bm{\Pi}=[\mathbf{U}]_{:,0:L-1}\times\sqrt{\bm{e}}$. In the following analysis, $\bm{\eta}$ is estimated, instead of $\bm{\theta}$.
III-A2 Estimation of approximate PN vector
By approximating the PN vector $\bm{\theta}$, the objective function on the right-hand side (RHS) of (8) can be rewritten as
$\mathcal{L}(\bm{\lambda})$ $=-\log p(\bm{y}\mid\bm{\lambda})-\log p(\bm{\eta})$
$\propto\frac{1}{\sigma^{2}}\|\bm{y}-\bm{\mu}\|^{2}+\frac{1}{2}\bm{\eta}^{\top}\bm{\eta},$ (10)
where $\bm{\mu}\triangleq\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\bm{h}$ stands for the noise-free received signal at the UE.
Owing to the typically small PN variance of practical oscillators, the elements of the PN matrix $\bm{\Lambda}_{\bm{\theta}}$ can be approximated by employing its first-order Taylor expansion, resulting in $e^{j\bm{\theta}[n]}\approx 1+j\bm{\theta}[n]$. Consequently, we can approximate $\bm{\Lambda}_{\bm{\theta}}$ as $\bm{\Lambda}_{\bm{\theta}}\approx\mathbf{I}_{N}+j{\rm Diag}(\bm{\Pi\eta})$. With this approximation, the objective function in (III-A2) can be rewritten as
$\mathcal{L}_{\bm{\eta}}^{[q+1]}$ $\approx\frac{1}{\sigma^{2}}(\bar{\bm{y}}-\mathbf{Q}\bm{\eta})^{\rm H}(\bar{\bm{y}}-\mathbf{Q}\bm{\eta})+\frac{1}{2}\bm{\eta}^{\top}\bm{\eta}$
$=\frac{1}{\sigma^{2}}\left(\bar{\bm{y}}^{\rm H}\bar{\bm{y}}-2\Re\left\{\bar{\bm{y}}^{\rm H}\mathbf{Q}\right\}\bm{\eta}+\bm{\eta}^{\top}\Re\left\{\mathbf{Q}^{\rm H}\mathbf{Q}\right\}\bm{\eta}\right)$
$\quad+\frac{1}{2}\bm{\eta}^{\top}\bm{\eta},$ (11)
where $\bar{\bm{y}}\triangleq\bm{y}-\sqrt{P}\hat{\bm{\Lambda}}_{\phi}^{[q]}\mathbf{S}^{\top}\hat{\bm{h}}^{[q]}$, and $\mathbf{Q}\triangleq j\sqrt{P}{\rm Diag}(\hat{\bm{\Lambda}}_{\phi}^{[q]}\mathbf{S}^{\top}\hat{\bm{h}}^{[q]})\bm{\Pi}$. The superscript ``$^{q}"$ indicates the $q$-th iteration.
Set the partial derivative of (III-A2) regarding $\bm{\eta}$ to zero, i.e.,
$\frac{\partial\mathcal{L}^{[q+1]}_{\bm{\eta}}}{\partial\bm{\eta}}=-\frac{2}{\sigma^{2}}\Re\left\{\mathbf{Q}^{\rm H}\bar{\bm{y}}\right\}+\frac{2}{\sigma^{2}}\Re\left\{\mathbf{Q}^{\rm H}\mathbf{Q}\right\}\bm{\eta}+\bm{\eta}=\bm{0}_{L\times 1},$ (12)
we can solve
$\hat{\bm{\eta}}^{[q+1]}=\left[\Re\left\{\mathbf{Q}^{\rm H}\mathbf{Q}\right\}+\frac{1}{2}\sigma^{2}\mathbf{I}_{L}\right]^{-1}\Re\left\{\mathbf{Q}^{\rm H}\bar{\bm{y}}\right\}.$ (13)
III-B CFO Estimation
To estimate the CFO, we solve the following unconstrained problem:
$\hat{\phi}\propto\arg\min_{\phi}\mathcal{L}_{\phi}\propto\arg\min_{\phi}\|\bm{y}-\bm{\mu}\|^{2}.$ (14)
To enhance the tractability of (14) and derive a closed-form solution, the first-order Taylor approximation is employed. The $n$-th term on the diagonal of $\bm{\Lambda}_{\phi}$ can be approximated as
$\left[\bm{\Lambda}_{\phi}\right]_{n,n}=e^{j2\pi\frac{n}{N}\phi}\approx e^{j2\pi\frac{n}{N}\hat{\phi}}+\left(\phi-\hat{\phi}\right)\frac{j2\pi}{N}e^{j2\pi\frac{n}{N}\hat{\phi}}.$ (15)
Hence, $\hat{\bm{\Lambda}}_{\phi}^{[q+1]}$ can be approximated as
$\hat{\bm{\Lambda}}_{\phi}^{[q+1]}\approx\hat{\bm{\Lambda}}_{\phi}^{[q]}+\left(\phi-\hat{\phi}^{[q]}\right)\tilde{\bm{\Lambda}}_{\phi}^{[q]},$ (16)
where $\left[\tilde{\bm{\Lambda}}_{\phi}^{[q]}\right]_{n,n}=\frac{j2\pi n}{N}e^{\frac{j2\pi n\hat{\phi}^{[q]}}{N}}$.
With (16), $\mathcal{L}_{\phi}$ in (14) can be rewritten as
$\mathcal{L}_{\phi}^{[q+1]}\approx\frac{1}{\sigma^{2}}\|{\bm{y}-\left[\hat{\bm{\Lambda}}_{\phi}^{[q]}+\left(\phi-\hat{\phi}^{[q]}\right)\tilde{\bm{\Lambda}}_{\phi}^{[q]}\right]\bm{d}}\|^{2},$ (17)
where $\bm{d}\triangleq\sqrt{P}\hat{\bm{\Lambda}}_{\bm{\theta}}^{[q+1]}\mathbf{S}^{\top}\hat{\bm{h}}^{[q]}$.
Upon setting the partial derivative of (17) with regard to $\phi$ to zero and solving it for $\phi$, we acquire the closed-form solution of $\hat{\phi}$ at the $(q+1)$-th iteration as
$\hat{\phi}^{[q+1]}=\hat{\phi}^{[q]}+\frac{\Re\left\{\left(\bm{y}-\hat{\bm{\Lambda}}_{\phi}^{[q]}\bm{d}\right)^{\rm H}\tilde{\bm{\Lambda}}_{\phi}^{[q]}\bm{d}\right\}}{\bm{d}^{\rm H}\tilde{\bm{\Lambda}}_{\phi}^{[q]{\rm H}}\tilde{\bm{\Lambda}}_{\phi}^{[q]}\bm{d}}.$ (18)
III-C Position Estimation
For position estimation, we employ the GD algorithm. Due to the same structures in the respective derivatives of the objective function with regard to the three position parameters $d_{Ru}$, $\phi_{az}$, and $\phi_{el}$, we use the symbol $\xi$ to represent any one of them. The partial derivative of the objective function with regard to $\xi$ is given by
$\frac{\partial\mathcal{L}_{\xi}}{\partial\xi}=\frac{2}{\sigma^{2}}\Re\left\{\left(P\bm{h}^{\rm H}\mathbf{S}^{*}\mathbf{S}^{\top}-\sqrt{P}\bm{y}^{\rm H}\bm{\Lambda}_{\bm{\theta}}\bm{\Lambda}_{\phi}\mathbf{S}^{\top}\right)\frac{\partial\bm{h}}{\partial\xi}\right\},$ (19)
where $\frac{\partial\bm{h}}{\partial\xi}$ is calculated by
$\frac{\partial\bm{h}}{\partial\xi}[n]$ $=\frac{\partial\bm{h}[n]}{\partial\xi}=-\sum_{r=1}^{N_{R}}\sqrt{\rho_{aR,n,r}}\sqrt{\rho_{Ru,n,r}}\bm{w}[r]$
×$\times e^{-j2\pi\frac{nB}{N}(\tau_{ar}+\tau_{ru})}\left(\frac{1}{d_{ru}}+j2\pi\frac{nB}{Nc}\right)\frac{\partial d_{ru}}{\partial\xi},$ (20)
with $\frac{\partial d_{ru}}{\partial\xi}$, $\xi=d_{Ru},\phi_{az},\phi_{el}$, being
$\frac{\partial d_{ru}}{\partial d_{Ru}}$ $=\frac{1}{d_{Ru}}\left[(d_{Ru}\sin\phi_{el}\cos\phi_{az}-x_{r})\sin\phi_{el}\cos\phi_{az}\right.$
$\left.\quad+(d_{Ru}\sin\phi_{el}\sin\phi_{az}-y_{r})\sin\phi_{el}\sin\phi_{az}\right.$
$\left.\quad+(d_{Ru}\cos\phi_{el}-z_{r})\cos\phi_{el}\right],$ (21)
$\frac{\partial d_{ru}}{\partial\phi_{az}}$ $=\frac{1}{d_{Ru}}\left[(d_{Ru}\sin\phi_{el}\cos\phi_{az}-x_{r})(-\sin\phi_{az}d_{Ru}\sin\phi_{el})\right.$
$\left.\quad+(d_{Ru}\sin\phi_{el}\sin\phi_{az}-y_{r})d_{Ru}\sin\phi_{el}\cos\phi_{az}\right],$ (22)
and
$\frac{\partial d_{ru}}{\partial\phi_{el}}$ $=\frac{1}{d_{Ru}}\left[(d_{Ru}\sin\phi_{el}\cos\phi_{az}-x_{r})d_{Ru}\cos\phi_{az}\cos\phi_{el}\right.$
$\left.\quad+(d_{Ru}\sin\phi_{el}\sin\phi_{az}-y_{r})d_{Ru}\sin\phi_{az}\cos\phi_{el}\right.$
$\left.\quad+(d_{Ru}\cos\phi_{el}-z_{r})(-d_{Ru}\sin\phi_{el})\right].$ (23)
III-D Overall Algorithm
The proposed joint estimation of the CFO, PN and UE position is summarized in Algorithm 1. Specifically, we initialize the parameters, including the initial estimation of the normalized CFO $\hat{\phi}$ and UE position $\hat{\bm{p}}_{u}$, the step size, $k_{p}$, the convergence accuracy for outer and inner loops, $\epsilon_{o}$ and $\epsilon_{i}$, and the maximum iteration number for the outer and inner loops, $N_{o}$ and $N_{i}$. After transforming the coordinates into the polar coordinates, we estimate the PN, CFO, and UE position alternately. Closed-form expressions are utilized to update the PN and CFO estimates in each iteration, while the GD algorithm is employed to estimate the UE position. This repeats until either the objective function $\mathcal{L}$ converges or the maximum number of iterations is reached. It is noteworthy that this AO-based algorithm takes the RIS phase shifts, i.e., $\bm{w}$, as input, which can affect the algorithm performance significantly. The configuration of the RIS phase shifts, $\bm{w}$, will be discussed in Section V.
In Algorithm 1, the objective function is expected to be non-increasing throughout iterations when updating the CFO and PN parameters because the approximate functions in (III-A2) and (17) show convexity with respect to $\bm{\eta}$ and $\phi$, respectively. In Steps 9, 10 and 11, with the UE position estimated in the previous iteration as input, the GD algorithm can ensure a lower objective function. Since the posterior probability is no larger than $1$, the negative logarithmic function $\mathcal{L}(\bm{\lambda})$ is non-negative and therefore lower bounded. Algorithm 1 converges.
The computational complexity for updating $\bm{\eta}$, $\bm{\theta}$, $\phi$, and position parameters $d_{Rt}$, $\phi_{az}$, and $\phi_{el}$ in Algorithm 1 are $\mathcal{O}(L^{3}+N^{2}+NL^{2})$, $\mathcal{O}(NL)$, $\mathcal{O}(N^{2})$, and $\mathcal{O}(\max\left\{N^{3},N_{R}\right\})$, respectively. Suppose the inner GD algorithm for updating the UE position converges after $K_{i}$ iterations. Consequently, the computational complexity of Algorithm 1 is $\mathcal{O}(\max\left\{L^{3}+N^{2}+NL^{2},K_{i}N^{3},K_{i}N_{R}\right\})$ per iteration.
Algorithm 1 Proposed iterative joint estimation algorithm.
1:  Input: $\bm{w}$;
2:  Initialization: $\hat{\phi}$, $\hat{\bm{p}}_{u}$, $k_{p}$, $\epsilon_{o}$, $\epsilon_{i}$, $N_{o}$, $N_{i}$;
3:  Calculate polar coordinate of UE using (1);
4:  repeat
5:     Given fixed $\hat{\bm{p}}_{u}$, update $\hat{\bm{\eta}}$ with $\hat{\phi}$ using (13);
6:     Update $\hat{\bm{\theta}}$ using (9);
7:     Given fixed $\hat{\bm{p}}_{u}$, update $\hat{\phi}$ with $\hat{\bm{\theta}}$ using (18);
8:     repeat
9:        $\hat{d}_{Ru}\leftarrow\hat{d}_{Ru}-\frac{\partial\mathcal{L}_{\xi}}{\partial d_{Ru}}k_{p}$;
10:        $\hat{\phi}_{az}\leftarrow\hat{\phi}_{az}-\frac{\partial\mathcal{L}_{\xi}}{\partial\phi_{az}}k_{p}$;
11:        $\hat{\phi}_{el}\leftarrow\hat{\phi}_{el}-\frac{\partial\mathcal{L}_{\xi}}{\partial\phi_{el}}k_{p}$;
12:     until $\mathcal{L}^{[p]}-\mathcal{L}^{[p+1]}\leq\epsilon_{i}$ after the $p$-th iteration or $p\textgreater N_{i}$.
13:  until $\mathcal{L}^{[q]}-\mathcal{L}^{[q+1]}\leq\epsilon_{o}$ after the $q$-th iteration or $q\textgreater N_{o}$.
14:  Calculate Cartesian coordinate of UE using (2);
15:  Output: $\hat{\phi}$, $\hat{\bm{\theta}}$, $\hat{\bm{p}}_{u}$.
III-E Performance Metrics
We specify the performance metrics for all estimation parameters in $\bm{\lambda}$ for the sake of evaluating Algorithm 1. For the position estimation, we take root mean square error (RMSE) as the performance index. For any testing UE position $\bm{p}_{u}$, ${\rm RMSE}_{u}$ can be expressed as
${\rm RMSE}_{u}=\sqrt{\frac{1}{M_{e}}\sum_{m=1}^{M_{e}}\|\hat{\bm{p}}_{u,m}-\bm{p}_{u}\|^{2}},$ (24)
where $M_{e}$ is the number of Monte Carlo tests, and $\hat{\bm{p}}_{u,m}$ denotes the estimation of the UE position in the $m$-th trial.
For the CFO and PN estimation, it can be straightforwardly observed from (6) that the received signal $\bm{y}$ does not change under a common phase rotation $\epsilon$, i.e.,
$\hat{\phi}\rightarrow\phi-\epsilon\quad\text{and }\quad\hat{\bm{\theta}}\rightarrow\bm{\theta}+\bm{\epsilon},$ (25)
where $\bm{\epsilon}[n]=\frac{2\pi n\epsilon}{N}$. Here, (25) indicates that there exists ambiguity in estimating the CFO and PN. The ambiguity leads to difficulties in achieving high estimation accuracy for individual parameters of the CFO and PN. Nevertheless, a high-precision joint estimation of CFO and PN is achievable, which is beneficial for estimating the UE position and will be illustrated in section VI.
To evaluate the joint estimation performance, we utilize a joint MSE measure, which can be calculated as
${\rm MSE}_{\phi,\bm{\theta}}=\mathbb{E}_{\phi,\bm{\theta}}\left(\|\bm{\underline{\gamma}}-\hat{\bm{\underline{\gamma}}}\|^{2}\right),$ (26)
where $\bm{\underline{\gamma}}=\bm{\gamma}-\bm{\gamma}[0]\bm{1}$, $\hat{\bm{\underline{\gamma}}}=\hat{\bm{\gamma}}-\hat{\bm{\gamma}}[0]\bm{1}$, $\bm{\gamma}=[\bm{\gamma}[0],\bm{\gamma}[1],\cdots,\bm{\gamma}[N-1]]^{\top}$, $\bm{\gamma}[n]=\bm{\theta}[n]+\frac{2\pi n\phi}{N}$, $\hat{\bm{\gamma}}=[\hat{\bm{\gamma}}[0],\hat{\bm{\gamma}}[1],\cdots,\hat{\bm{\gamma}}[N-1]]^{\top}$, and $\hat{\bm{\gamma}}[n]=\hat{\bm{\theta}}[n]+\frac{2\pi n\hat{\phi}}{N}$. Using this measure, the performance of the joint CFO-PN estimation rather than each of the individual variables alone is considered.
IV Hybrid Cramér-Rao Lower Bound Analysis
The HCRLB for the joint estimation of the CFO, PN, and UE position is derived in this section. This helps quantify the performance lower bound of the proposed AO-based estimation algorithm, and also serves as the objective function for optimizing the RIS phase shifts, $\bm{w}$, in the next section. The HCRLB for the estimation of $\bm{\lambda}$ is given by [34, Eq. (4.610)]
$\mathbb{E}_{\bm{y},\bm{\theta}\mid\phi,d_{Ru},\phi_{az},\phi_{el}}\left[(\bm{\lambda}-\hat{\bm{\lambda}})(\bm{\lambda}-\hat{\bm{\lambda}})^{\top}\right]\succeq\mathbf{B}^{\rm-1},$ (27)
where $\mathbf{B}$ is the Bayesian information matrix (BIM), which can be given by [34, Eq. (4.611)]
$\mathbf{B}=\mathbb{E}_{\bm{\theta}}\left[\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})\right]+\mathbb{E}_{\bm{\theta}}\left[-\frac{\partial}{\partial\bm{\lambda}}\left(\frac{\partial\log p(\bm{\theta})}{\partial\bm{\lambda}}\right)^{\top}\right],$ (28)
where $\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})$ stands for the Fisher’s information matrix (FIM) for $\bm{\lambda}$ under the observation of received signal $\bm{y}$ at the UE, and $\mathbb{E}_{\bm{\theta}}\left[-\frac{\partial}{\partial\bm{\lambda}}\left(\frac{\partial\log p(\bm{\theta})}{\partial\bm{\lambda}}\right)^{\top}\right]$ is the prior information matrix. Interested readers may refer to [34] for more details.
IV-A Derivation of $\mathbb{E}_{\bm{\theta}}\left[\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})\right]$
To derive $\mathbb{E}_{\bm{\theta}}\left[\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})\right]$ in (28), we first analyze the FIM for $\bm{\lambda}$, as given by
$\left[\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})\right]_{i,j}=\frac{2}{\sigma^{2}}\Re\left\{\frac{\partial\bm{\mu}{\rm{}^{\rm H}}}{\partial\lambda_{i}}\frac{\partial\bm{\mu}}{\partial\lambda_{j}}\right\}.$ (29)
To calculate (29), the first-order derivative of $\bm{\mu}$ with respect to each element in vector $\bm{\lambda}$ is derived as follows.
$\frac{\partial\bm{\mu}}{\partial\phi}$ $=\sqrt{P}\bm{\Lambda}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\bm{h};$ (30a)
$\frac{\partial\bm{\mu}}{\partial\bm{\theta}[n]}$ $=\sqrt{P}{\rm Diag}(\bm{a}[n])\bm{\Lambda}_{\phi}\mathbf{S}^{\top}\bm{h};$ (30b)
$\frac{\partial\bm{\mu}}{\partial d_{Ru}}$ $=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial d_{Ru}};$ (30c)
$\frac{\partial\bm{\mu}}{\partial\phi_{az}}$ $=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial\phi_{az}};$ (30d)
$\frac{\partial\bm{\mu}}{\partial\phi_{el}}$ $=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial\phi_{el}},$ (30e)
where $\bm{\Lambda}$ is a diagnal matrix, $\left[\bm{\Lambda}\right]_{n,n}=\frac{j2\pi n}{N}$, and $\bm{a}[n]=[\bm{0}_{1\times n},je^{j\bm{\theta}[n]},\bm{0}_{1\times(N-n-1)}]^{\top}$.
It is not difficult to see that in (30), only $\bm{\Lambda}_{\phi}$, $\bm{\Lambda}_{\bm{\theta}}$ and ${\rm Diag}(\bm{a}[n])$ are related to $\phi$ and $\bm{\theta}$, which are all diagonal matrices. Additionally, both $\phi$ and $\bm{\theta}$ appear in the exponent of diagonal elements in these matrices. Hence, they can be suppressed when calculating $\frac{\partial\bm{\mu}{\rm{}^{\rm H}}}{\partial\lambda_{i}}\frac{\partial\bm{\mu}}{\partial\lambda_{j}}$ in (29). In other words, the calculated FIM does not rely on the parameter $\bm{\theta}$, i.e.,
$\mathbb{E}_{\bm{\theta}}\left[\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})\right]=\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda}).$ (31)
Take $\frac{\partial\bm{\mu}^{\rm H}}{\partial d_{Ru}}\frac{\partial\bm{\mu}}{\partial d_{Ru}}$ as an example. We have
$\frac{\partial\bm{\mu}^{\rm H}}{\partial d_{Ru}}\frac{\partial\bm{\mu}}{\partial d_{Ru}}$ $=\sqrt{P}\frac{\partial\bm{h}^{\rm H}}{\partial d_{Ru}}\mathbf{S}^{*}\bm{\Lambda}_{\bm{\theta}}^{\rm{H}}\bm{\Lambda}_{\phi}^{\rm H}\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial d_{Ru}}$ (32)
$=P\frac{\partial\bm{h}^{\rm H}}{\partial d_{Ru}}\mathbf{S}^{*}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial d_{Ru}}$
which is unrelated to $\phi$ or $\bm{\theta}$, since $\bm{\Lambda}_{\phi}^{\rm H}\bm{\Lambda}_{\phi}=\bm{\Lambda}_{\bm{\theta}}^{\rm{H}}\bm{\Lambda}_{\bm{\theta}}=\mathbf{I}_{N}$. This can be extended to $\frac{\partial\bm{\mu}{\rm{}^{\rm H}}}{\partial\lambda_{i}}\frac{\partial\bm{\mu}}{\partial\lambda_{j}},\forall i,j$, and is omitted here.
IV-B Derivation of $\mathbb{E}_{\bm{\theta}}\left[-\frac{\partial}{\partial\bm{\lambda}}\left(\frac{\partial\log p(\bm{\theta})}{\partial\bm{\lambda}}\right)^{\top}\right]$
Since $\bm{\theta}\sim\mathcal{N}(\bm{0}_{N\times 1},\bm{\Psi})$, we can calculate $\log p(\bm{\theta})$ as $\log p(\bm{\theta})=-\frac{1}{2}\bm{\theta}^{\top}\bm{\Psi}^{-1}\bm{\theta}-\frac{N}{2}\log(2\pi)-\frac{1}{2}\log\det\bm{\Psi}$, and its partial derivative with respect to $\bm{\lambda}$ can be calculated by
$\frac{\partial\log p(\bm{\theta})}{\partial\bm{\lambda}}=\left[0,-(\bm{\Psi}^{-1}\bm{\theta})^{\top},0,0,0\right]^{\top}.$ (33)
Subsequently, $\mathbb{E}_{\bm{\theta}}\left[-\frac{\partial}{\partial\bm{\lambda}}\left(\frac{\partial\log p(\bm{\theta})}{\partial\bm{\lambda}}\right)^{\top}\right]$ can be expressed by
$\mathbb{E}_{\bm{\theta}}\left[-\frac{\partial}{\partial\bm{\lambda}}\left(\frac{\partial\log p(\bm{\theta})}{\partial\bm{\lambda}}\right)^{\top}\right]$ $=-\frac{\partial}{\partial\bm{\lambda}}\left(\frac{\partial\log p(\bm{\theta})}{\partial\bm{\lambda}}\right)^{\top}$
$=\mathrm{Blkdiag}(0,\bm{\Psi}^{-1},\bm{0}_{3\times 3}).$ (34)
IV-C Derivation of the Transformed HCRLB
So far, we have derived the expression for $\mathbf{B}$, which is an HCRLB matrix that gives analytical estimation accuracy lower bounds for the parameters in $\bm{\lambda}$. However, when assessing the proposed estimation algorithm, we adopt the RMSE and CFO-PN joint MSE metrics introduced in Section III-E. Hence, we define a new parameter vector $\bm{\alpha}=\bm{g}(\bm{\lambda})\triangleq\left[\bm{\underline{\gamma}},x_{u},y_{u},z_{u}\right]^{\top}$, corresponding to the two metrics. Then, $\mathbf{B}$ can be transformed into the $\bm{\alpha}$-related HCRLB matrix by [35]
$\mathbf{HCRLB}(\bm{p}_{u},\bm{w})=\bm{\Xi}(\bm{p}_{u})\mathbf{B}^{-1}(\bm{p}_{u},\bm{w})\bm{\Xi}^{\top}({\bm{p}_{u}}),$ (35)
where, for any UE position $\bm{p}_{u}$, $\mathbf{\Xi}(\bm{p}_{u})$ is the transition matrix defined as
$\bm{\Xi}(\bm{p}_{u})=\frac{\partial\bm{g}(\bm{\lambda})}{\partial\bm{\lambda}}={\rm Blkdiag}\left(\bm{\Xi}_{1},\bm{\Xi}_{2}(\bm{p}_{u})\right),$ (36)
with $\bm{\Xi}_{1}$ and $\bm{\Xi}_{2}(\bm{p}_{u})$ given by (37) and (38), respectively.
$\bm{\Xi}_{1}=\begin{bmatrix}\frac{\partial\bm{\underline{\gamma}}[0]}{\partial\phi}&\frac{\partial\bm{\underline{\gamma}}[0]}{\partial\bm{\theta}[0]}&\cdots&\frac{\partial\bm{\underline{\gamma}}[0]}{\partial\bm{\theta}[N-1]}\\ \vdots&\vdots&\ddots&\vdots\\ \frac{\partial\bm{\underline{\gamma}}[N-1]}{\partial\phi}&\frac{\partial\bm{\underline{\gamma}}[N-1]}{\partial\bm{\theta}[0]}&\cdots&\frac{\partial\bm{\underline{\gamma}}[N-1]}{\partial\bm{\theta}[N-1]}\end{bmatrix}=\begin{bmatrix}0&0&0&\cdots&0\\ \frac{2\pi}{N}&-1&1&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \frac{2\pi(N-1)}{N}&-1&0&\cdots&1\\ \end{bmatrix}\in\mathbb{R}^{N\times(N+1)},$ (37)
$\bm{\Xi}_{2}(\bm{p}_{u})=\begin{bmatrix}\frac{\partial x_{u}}{\partial d_{Ru}}&\frac{\partial x_{u}}{\partial\phi_{az}}&\frac{\partial x_{u}}{\partial\phi_{el}}\\ \frac{\partial y_{u}}{\partial d_{Ru}}&\frac{\partial y_{u}}{\partial\phi_{az}}&\frac{\partial y_{u}}{\partial\phi_{el}}\\ \frac{\partial z_{u}}{\partial d_{Ru}}&\frac{\partial z_{u}}{\partial\phi_{az}}&\frac{\partial z_{u}}{\partial\phi_{el}}\\ \end{bmatrix}=\begin{bmatrix}\sin\phi_{el}\cos\phi_{az}&-d_{Ru}\sin\phi_{el}\sin\phi_{az}&d_{Ru}\cos\phi_{az}\cos\phi_{el}\\ \sin\phi_{el}\sin\phi_{az}&d_{Ru}\sin\phi_{el}\cos\phi_{az}&d_{Ru}\sin\phi_{az}\cos\phi_{el}\\ \cos\phi_{el}&0&-d_{Ru}\sin\phi_{el}\end{bmatrix}\in\mathbb{R}^{3\times 3}.$ (38)
After obtaining $\mathbf{HCRLB}(\bm{p}_{u},\bm{w})$, the position error bound (${\rm PEB}$) for evaluating the UE position estimation and $\rm HCRLB_{CFO-PN}$ for evaluating the CFO-PN joint estimation can be given by
${\rm PEB}=\sqrt{{\rm tr}([\mathbf{HCRLB}(\bm{p}_{u},\bm{w})]_{N:N+2,N:N+2})};$ (39)
${\rm HCRLB_{CFO-PN}}={\rm tr}([\mathbf{HCRLB}(\bm{p}_{u},\bm{w})]_{0:N-1,0:N-1}).$ (40)
Both ${\rm PEB}$ and ${\rm HCRLB_{CFO-PN}}$ depend on the UE’s position, $\bm{p}_{u}$, and the RIS phase shifts, $\bm{w}$.
V RIS Configuration
In this section, we further configure the phase shifts $\bm{w}$ of the RIS to minimize the analytical lower bound of the localization accuracy, i.e., PEB. According to (39), the PEB depends on the UE position, which indicates that minimizing the PEB at a specific UE position requires knowledge of that position. However, this is not possible in practice. To address this issue and enhance the algorithm’s robustness, we take the average PEB within the area of interest (AOI) as the objective function. Specifically, we randomly sample $U$ positions within the AOI and calculate their average PEB. Accordingly, the problem of interest is formulated as
$\min_{\bm{w}}\quad\frac{1}{U}\sum_{u=1}^{U}{\rm PEB}(\bm{p}_{u},\bm{w})$
$s.t.\quad|\bm{w}[r]|=1,r=1,\cdots,N_{R}.$ (41)
Problem (V) is not convex because of the non-convexity of the objective function with respect to vector $\bm{w}$, as well as the unit-module constraint.
To convexify (V), we first define $\mathbf{W}\triangleq\bm{\bar{w}}\bm{\bar{w}}^{\rm H}$, where $\bm{\bar{w}}\triangleq\bm{w}^{*}$, which satisfies $\mathbf{W}\succcurlyeq 0$ and ${\rm rank}(\mathbf{W})=1$. Since the rank-1 constraint is non-convex, we employ the SDR method to relax this constraint. Consequently, the problem in (V) can be transformed into
$\min_{\mathbf{W}}\quad\frac{1}{U}\sum_{u=1}^{U}{\rm PEB}(\bm{p}_{u},\mathbf{W})$
$\begin{array}[]{ll}s.t.&\left[\mathbf{W}\right]_{r,r}=1,r=1,\cdots,N_{R},\\ &\mathbf{W}\succcurlyeq 0.\end{array}$ (44)
According to (35) and (39), the optimization problem in (V) remains non-convex because of the complex objective function. To transform (V) into the convex SDP form, for each calculated PEB at a particular position $\bm{p}_{u}$, we define an auxiliary matrix $\mathbf{Z}_{u}$ that satisfies
$\mathbf{Z}_{u}\succcurlyeq\tilde{\bm{\Xi}}(\bm{p}_{u})\mathbf{B}^{-1}(\bm{p}_{u},\mathbf{W})\tilde{\bm{\Xi}}^{\top}(\bm{p}_{u}),$ (45)
where $\tilde{\bm{\Xi}}(\bm{p}_{u})\triangleq\mathbf{T}\mathbf{\Xi}(\bm{p}_{u})$, and $\mathbf{T}\triangleq[\bm{0}_{N\times N}\quad\mathbf{I}_{3}]\in\mathbb{R}^{(N+3)\times(N+3)}$ can extract the PEB from the HCRLB. According to the property of the Schur complement and the positive semidefiniteness of $\mathbf{B}(\bm{p}_{u},\mathbf{W})$, (45) can be transformed to
$\begin{bmatrix}\mathbf{Z}_{u}&\tilde{\bm{\Xi}}(\bm{p}_{u})\\ \tilde{\bm{\Xi}}^{\top}(\bm{p}_{u})&\mathbf{B}(\bm{p}_{u},\mathbf{W})\end{bmatrix}\succcurlyeq 0.$ (46)
Thus, the problem in (V) can be converted to
$\min_{\mathbf{W},\mathbf{Z}_{1},\cdots,\mathbf{Z}_{U}}\quad\frac{1}{U}\sum_{u=1}^{U}{\rm tr}(\mathbf{Z}_{u})$
$\begin{array}[]{ll}s.t.&\left[\mathbf{W}\right]_{r,r}=1,r=1,\cdots,N_{R},\\ &\mathbf{W}\succcurlyeq 0,\\ &\displaystyle\frac{1}{U}\sum_{u=1}^{U}\begin{bmatrix}\mathbf{Z}_{u}&\tilde{\bm{\Xi}}(\bm{p}_{u})\\ \tilde{\bm{\Xi}}^{\top}(\bm{p}_{u})&\mathbf{B}(\bm{p}_{u},\mathbf{W})\end{bmatrix}\succcurlyeq 0,\end{array}$ (50)
where, according to the relationship in (45) and the calculation of PEB in (39), minimizing the objective function results in minimizing the average square of PEB, which also minimizes the initial objective function, the average PEB.
Proposition 1. $\mathbf{B}(\bm{p}_{u},\mathbf{W})$ is linear to $\mathbf{W}$.
Proof. According to (28), only $\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})$ depends on $\bm{w}$ in $\mathbf{B}$. Hence, we only need to prove the linearity of $\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})$ to $\mathbf{W}$, which can be rewritten as
$\left[\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})\right]_{i,j}=\frac{2}{\sigma^{2}}\sum_{n=0}^{N-1}\Re\left\{\frac{\partial\bm{\mu}[n]^{*}}{\partial\lambda_{i}}\frac{\partial\bm{\mu}[n]}{\partial\lambda_{j}}\right\}.$ (51)
To derive the explicit expressions of $\frac{\partial\bm{\mu}[n]^{*}}{\partial\lambda_{i}},\forall i$ and $\frac{\partial\bm{\mu}[n]}{\partial\lambda_{j}},\forall j$ with respect to $\bar{\bm{w}}$, we rewrite (6) as
$\bm{y}=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\mathbf{G}^{\top}\bm{w}+\bm{v},$ (52)
with $\mathbf{G}\in\mathbb{C}^{N_{R}\times N}$ and $\left[\mathbf{G}\right]_{r,n}=\sqrt{\rho_{aR,n,r}}\sqrt{\rho_{Ru,n,r}}e^{-j2\pi\frac{nB}{N}(\tau_{ar}+\tau_{ru})}$.
Correspondingly, $\bm{\mu}$ can be rewritten as $\bm{\mu}=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\mathbf{G}^{\top}\bm{\bar{w}}^{*}$. Furthermore, we can express each term of $\bm{\mu}$ and its conjugate as
$\bm{\mu}[n]=\bm{\mu}[n]^{\top}$ $=\sqrt{P}e^{j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}[\mathbf{S}^{\top}]_{n,:}\mathbf{G}^{\top}\bm{\bar{w}}^{*}$
$=\sqrt{P}e^{j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}\bm{\bar{w}}^{\rm H}\mathbf{G}[\mathbf{S}]_{:,n},$ (53)
and
$\bm{\mu}[n]^{*}=\sqrt{P}e^{-j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}[\mathbf{S}]_{:,n}^{\rm H}\mathbf{G}^{\rm H}\bm{\bar{w}}.$ (54)
The partial derivatives of $\bm{\mu}[n]$ and $\bm{\mu}[n]^{*}$ with regard to all parameters to be estimated are derived in Appendix A. It can be observed that $\bm{\bar{w}}$ is only in the last term of $\frac{\partial\bm{\mu}[n]^{*}}{\partial\lambda_{i}},\forall i$, while $\bm{\bar{w}}^{\rm H}$ appears as the first term of $\frac{\partial\bm{\mu}[n]}{\partial\lambda_{j}},\forall j$, for the vector parts. Hence, the product of $\frac{\partial\bm{\mu}[n]^{*}}{\partial\lambda_{i}},\forall i$ and $\frac{\partial\bm{\mu}[n]}{\partial\lambda_{j}},\forall j$ would lead to $\mathbf{W}$. As a result, as can be seen from (51), every term of $\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})$ is linear to $\mathbf{W}$ because the operations of taking the real parts, summation, and scalar multiplication preserve linearity. Hence, $\mathbf{B}(\bm{p}_{u},\mathbf{W})$ is linear to $\mathbf{W}$. ∎
According to Proposition 1, (V) is a convex SDP problem, and can be optimally solved using established convex optimization solvers, such as off-the-shelf CVX toolboxes [36]. Since the rank-1 constraint is relaxed via the SDR process, the obtained matrix $\mathbf{W}$ is likely to have a rank greater than 1. Therefore, decomposing it into the desired RIS phase shift vector $\bm{w}$ is a crucial subsequent problem. As done in [37], we first perform EVD on the optimal $\mathbf{W}$, then take the largest eigenvector and multiply it with the square root of the largest eigenvalue and normalize the vector, obtaining $\bm{\bar{w}}$. Finally, we calculate the RIS phase shifts by $\bm{w}=\bm{\bar{w}}^{*}$. The decomposition process can also be accomplished using the Gaussian randomization method, as done in [38].
Being an SDP, the problem in (V) can be solved using the interior point method in the CVX toolboxes. Thus, the convergence of the proposed RIS configuration is guaranteed. The computational complexity of the proposed RIS configuration is primarily influenced by the interior point method employed to solve (V). The worst-case complexity of the interior point method is $\mathcal{O}(\max\left\{N_{con},N_{var}\right\}^{4}\sqrt{N_{v}})$, where $N_{con}$ and $N_{var}$ are the constraint and variable numbers, respectively [39]. For the problem in (V), ${N_{con}}=N_{R}+2$ and ${N_{var}}=N_{R}^{2}+U(N+3)^{2}$. Suppose that the convergence accuracy is $\epsilon$. The complexity of solving (V) is $\mathcal{O}((N_{R}^{2}+UN^{2})^{4.5}\log\frac{1}{\epsilon})$.
VI Numerical Evaluation
This section elaborates on the comprehensive simulation results conducted to assess the proposed CFO, PN and position estimation and the RIS configuration algorithms. We adopt a 3D Cartesian coordinate system, where the AN is located at $\bm{p}_{a}=[\rm 2\>m,-2\>m,1\>m]^{\top}$, and the reference point of the RIS is located at $\bm{0}_{3\times 1}{\rm\>m}$. The RIS is placed on the $yz$ plane, without loss of generality. The AOI is set to be a cube centered at $[\rm 2\>m,2\>m,0\>m]^{\top}$ with the side length of $\rm 1\>m$. The light speed is $c=3\times 10^{8}\>{\rm m/s}$, and the noise power is $\sigma^{2}=-109\;{\rm dBm}$. $N=32$ subcarriers are used in an OFDM symbol, the central frequency is $f_{c}=2.8{\rm\>GHz}$, and the bandwidth is $B=100{\rm\>MHz}$. The antenna gains $G_{a}$ and $G_{t}$ are both set to $1$. Unless otherwise specified, $N_{R}=81$ and $\sigma_{\Delta}^{2}=10^{-3}\;{\rm rad^{2}}$ are used in the simulations.
The length of the approximate PN vector $\bm{\eta}$ is $L=16$. The normalized CFO $\phi$ is uniformly drawn from $[-0.15,0.15]$, corresponding to a CFO of up to $15\;{\rm MHz}$. The CFO can be caused by frequency mismatches (resulting from severe hardware impairments) between the transmitter (the ANs) and receiver (the UE) when the UE is stationary. [^1] The CFO can also be a result of the Doppler shift, in which case the normalized CFO is expected to be within a much smaller range, e.g., $[-9.96\times 10^{-5},9.96\times 10^{-5}]$ when the UE moves at the speed of $\rm 120\;km/h$. The proposed localization algorithm can directly apply to this case due to the direct mapping between the Doppler shift and the normalized CFO.
We gauge the performance of the proposed algorithms in comparison with the state-of-the-art approaches developed in [27], where a 2D near-field localization problem was solved with the assistance of a RIS. An MLE-based algorithm was designed to estimate the UE position, and an SDR-based optimization algorithm was proposed to minimize the CRLB. Note that the CFO and PN were not considered in [27]. The RMSE and the CFO-PN joint MSE are utilized to assess the performance of the localization and joint CFO and PN estimation; see (24) and (26).
VI-A Convergence of Algorithm 1
(a) Convergence of the GD algorithm for estimating UE position.
(b) Convergence of the overall algorithm.
Figure 3: The convergence of the proposed Algorithm 1, where the transmit power is $P=-10{\rm\;dBm}$.
First, we illustrate the convergence of Algorithm 1. Fig. 3 illustrates the case when the transmit power is $-10{\rm\>dBm}$, which corresponds to a receive signal-to-noise ratio (SNR) of approximately $17.5{\rm\>dB}$. Specifically, Fig. 3(a) depicts the change of the objective function when operating the GD algorithm to estimate the UE position, and different curves correspond to the GD algorithm in different rounds of iteration. Fig. 3(b) focuses on the convergence of the overall algorithm, and different points on the curves represent the objective function’s values after different rounds of iteration. In general, the GD algorithm typically converges within 2,000 iterations in this setup, which converges increasingly earlier within an iteration as the overall algorithm runs. The overall algorithm generally converges within 55 iterations.
TABLE II: Average number of iterations for Algorithm 1
Transmit Power (dBm) $-20$ $-10$ 00 $10$
Avg. No. of Iterations 19.45 50.59 84.31 134.95
Further, Table II shows the average number of iterations under different transmit powers for Algorithm 1. In this simulation, 100 tests are conducted, and $N_{o}$ and $N_{i}$ are set to be sufficiently large. The results indicate that the required iteration number increases as the transmit power increases.
VI-B Performance of RIS Configuration
(a) Random RIS phase shift
(b) Approach developed in [27]
(c) Proposed optimization algorithm
Figure 4: The PEB within AOI for various RIS configuration strategies.
Next, our attention is directed towards evaluating the performance of the proposed RIS configuration algorithm. Fig. 4 illustrates the PEB near the AOI across various RIS phase shift design schemes, where the black squares represent the boundary of AOI. More precisely, we set the transmit power $P$ to $-20{\rm\>dBm}$, corresponding to a receive SNR of approximately $7.5{\rm\>dB}$. Without loss of generality, the selected positions in the figures are all located in the $xy$-plane.
Fig. 4 indicates that the proposed optimization algorithm improves the PEB by approximately two orders of magnitude by comparison with the random RIS phase shifts, and by approximately one order of magnitude compared to the situation where there are no hardware impairments [27]. This fully demonstrates the significance of considering the CFO and PN in the localization of the UE, as well as the effectiveness, progressiveness, and superiority of the proposed RIS configuration algorithm. Additionally, observed from Fig. 4, the optimized RIS phase shifts not only improve the localization accuracy within the AOI but also significantly improve localization performance in its surrounding areas.
VI-C Estimation Accuracy
Figure 5: The CDF of RMSE and PEB with different $N_{R}$ and transmit powers.
Then, we concentrate on the CFO, PN, and UE position estimation accuracy. We first illustrate the overall localization accuracy. Fig. 5 illustrates the cumulative distribution functions (CDFs) of RMSE and PEB for 500 random testing positions, under different transmit powers and different numbers of RIS elements. It is observed that increasing the RIS size and the transmit power results in significant enhancement of the overall localization accuracy. Furthermore, when the RIS size is $N_{R}=49$, the CDF of the RMSE deviates significantly from the CDF of the PEB, and does not achieve the theoretical localization accuracy. However, when the RIS size is increased to $N_{R}=121$, the CDF of the RMSE curve aligns more closely with the CDF of the PEB. This phenomenon can also be observed in Fig. 8.
Next, we select a fixed testing position with the coordinates $\rm[1.50\>m,2.15\>m,0.45\>m]^{\top}$ to investigate the impact of the transmit power $P$, the RIS size $N_{R}$, and the PN variance $\sigma_{\Delta}^{2}$ on the CFO, PN, and UE position estimation accuracy.
Figure 6: The RMSE of $\bm{p}_{u}$ estimation under random and optimized RIS phase shifts.
Figure 7: ${\rm MSE}_{\phi,\bm{\theta}}$ under random and optimized RIS phase shifts.
In Figs. 6 and 7, we assess the effect of the optimized RIS phase shifts on the performance of estimation, compared to the random setup. From Fig. 6, the localization performance with random RIS phase shifts does not experience a significant decline as the transmit power increases. At the other extreme, the optimized RIS phase shifts provide an improvement of approximately two orders of magnitude in localization accuracy. This demonstrates the crucial role of optimizing the RIS phase shifts in localization and underscores the effectiveness of the proposed RIS configuration optimization algorithm. On the other hand, Fig. 7 shows that compared to the optimized RIS phase shifts, using random RIS phase shifts yields even lower joint MSE of CFO and PN, although the curve does not experience any significant decline as the transmit power increases in this case. This occurs because we extract the PEB part of HCRLB and set it as the RIS optimization objective. Thus, the estimation performance of the CFO and PN parameters is ignored.
Figure 8: The RMSE of $\bm{p}_{u}$ estimation at $N_{R}=49$, $N_{R}=81$, and $N_{R}=121$.
Figure 9: ${\rm MSE}_{\phi,\bm{\theta}}$ at $N_{R}=49$, $N_{R}=81$, and $N_{R}=121$.
Figs. 8 and 9 depict the curves of the RMSE for UE localization and MSE for the estimation of $\underline{\bm{\gamma}}$, respectively, under different transmit powers, where $N_{R}$ is 49, 81, or 121. It is observed that increasing the transmit power and/or the RIS element number results in gains in all parameters to be estimated. However, as the transmit SNR increases, the RMSE and MSE do not decrease as sharply as HCRLB because of the approximate PN vector, as well as the Taylor series approximation considered when deriving closed-form expressions for the CFO and PN.
Figure 10: ${\rm RMSE}_{u}$ at $\sigma_{\Delta}^{2}=10^{-3}{\rm\>rad^{2}}$ and $\sigma_{\Delta}^{2}=10^{-4}{\rm\>rad^{2}}$.
Figure 11: ${\rm MSE}_{\phi,\bm{\theta}}$ at $\sigma_{\Delta}^{2}=10^{-3}{\rm\>rad^{2}}$ and $\sigma_{\Delta}^{2}=10^{-4}{\rm\>rad^{2}}$.
Figs. 10 and 11 plot the RMSE for localization and MSE for estimation of $\underline{\bm{\gamma}}$, respectively, when $N_{R}$ is set to 81, where the PN variance is $\sigma_{\Delta}^{2}=10^{-3}{\rm\>rad^{2}}$ or $\sigma_{\Delta}^{2}=10^{-4}{\rm\>rad^{2}}$. The results demonstrate that decreasing the PN covariance leads to significant improvements in the estimation performance of the CFO, PN, and UE position. Another important finding is that a smaller PN variance would make the RMSE and MSE curves more closely aligned with the HCRLB curves because of the more accurately approximated parameters when updating the CFO and PN parameters.
Figure 12: Benchmarking the performance of the proposed algorithms against the state-of-the-art approach in [27].
Last but not least, we compare the proposed overall algorithm of the CFO, PN, and UE position estimation and the RIS configuration, with the benchmark algorithm developed in [27]. It is shown in Fig. 12 that, after optimization, the theoretical localization accuracy of our proposed algorithm outperforms the benchmark, which confirms the conclusion drawn in Fig. 4(b). As for the localization RMSE, the benchmark shows a large gap with its corresponding PEB, and its performance shows no demonstrable correlation with the transmit power because the estimation algorithm totally overlooked the CFO and PN. In comparison, the proposed algorithm significantly improves the localization accuracy by about one order of magnitude under relatively low transmit powers of below $-10{\rm\>dBm}$, and by two orders of magnitude under high transmit powers of over $10{\rm\>dBm}$.
VII Conclusion
This paper investigates the optimization of RIS phase shifts and the joint estimation of CFO, PN, and UE position under imperfect hardware conditions for an OFDM system in a near-field static localization scenario. An AO-based, iterative joint estimation algorithm based on the MAP criterion and GD algorithm has been proposed. Extensive simulations have demonstrated that by optimizing the RIS phase shifts, the theoretical localization accuracy has been improved by two orders of magnitude. Moreover, the achievable localization accuracy can be lower than $\rm 10^{-2}\>m$. These findings corroborate the effectiveness of the proposed algorithms. To further improve the estimation performance and support a larger AOI, the optimization algorithm used in this paper requires averaging a large number of PEBs related to random positions with high computational complexity. Hence, it becomes imperative to utilize an optimization algorithm with reduced complexity in this scenario. This will be our future work.
Appendix A Derivation of partial derivatives of $\bm{\mu}[n]$ and $\bm{\mu}[n]^{*}$
The partial derivatives of $\bm{\mu}[n]$ with regard to all parameters to be estimated are derived as follows
$\frac{\partial\bm{\mu}[n]}{\partial\phi}$ $=\sqrt{P}j\frac{2\pi n}{N}e^{j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}\bm{\bar{w}}^{\rm H}\mathbf{G}[\mathbf{S}]_{:,n},$ (55a)
$\frac{\partial\bm{\mu}[n]}{\partial\bm{\theta}[n]}$ $=\sqrt{P}je^{j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}\bm{\bar{w}}^{\rm H}\mathbf{G}[\mathbf{S}]_{:,n},$ (55b)
$\frac{\partial\bm{\mu}[n]}{\partial\bm{\theta}[m]}$ $=0,m\neq n,$ (55c)
$\frac{\partial\bm{\mu}[n]}{\partial d_{Ru}}$ $=\sqrt{P}e^{j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}\bm{\bar{w}}^{\rm H}\frac{\partial\mathbf{G}}{\partial d_{Ru}}[\mathbf{S}]_{:,n},$ (55d)
$\frac{\partial\bm{\mu}[n]}{\partial\phi_{az}}$ $=\sqrt{P}e^{j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}\bm{\bar{w}}^{\rm H}\frac{\partial\mathbf{G}}{\partial\phi_{az}}[\mathbf{S}]_{:,n},$ (55e)
$\frac{\partial\bm{\mu}[n]}{\partial\phi_{el}}$ $=\sqrt{P}e^{j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}\bm{\bar{w}}^{\rm H}\frac{\partial\mathbf{G}}{\partial\phi_{el}}[\mathbf{S}]_{:,n},$ (55f)
and the partial derivatives of $\bm{\mu}[n]^{*}$ with regard to all parameters to be estimated are derived as follows
$\frac{\partial\bm{\mu}[n]^{*}}{\partial\phi}$ $=-j\sqrt{P}\frac{2\pi n}{N}e^{-j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}[\mathbf{S}]_{:,n}^{\rm H}\mathbf{G}^{\rm H}\bm{\bar{w}},$ (56a)
$\frac{\partial\bm{\mu}[n]^{*}}{\partial\bm{\theta}[n]}$ $=-j\sqrt{P}e^{-j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}[\mathbf{S}]_{:,n}^{\rm H}\mathbf{G}^{\rm H}\bm{\bar{w}},$ (56b)
$\frac{\partial\bm{\mu}[n]^{*}}{\partial\bm{\theta}[m]}$ $=0,m\neq n,$ (56c)
$\frac{\partial\bm{\mu}[n]^{*}}{\partial d_{Ru}}$ $=\sqrt{P}e^{-j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}[\mathbf{S}]_{:,n}^{\rm H}\left(\frac{\partial\mathbf{G}}{\partial d_{Ru}}\right)^{\rm H}\bm{\bar{w}},$ (56d)
$\frac{\partial\bm{\mu}[n]^{*}}{\partial\phi_{az}}$ $=\sqrt{P}e^{-j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}[\mathbf{S}]_{:,n}^{\rm H}\left(\frac{\partial\mathbf{G}}{\partial\phi_{az}}\right)^{\rm H}\bm{\bar{w}},$ (56e)
$\frac{\partial\bm{\mu}[n]^{*}}{\partial\phi_{el}}$ $=\sqrt{P}e^{-j(\bm{\theta}[n]+2\pi\frac{n}{N}\phi)}[\mathbf{S}]_{:,n}^{\rm H}\left(\frac{\partial\mathbf{G}}{\partial\phi_{el}}\right)^{\rm H}\bm{\bar{w}}.$ (56f)